### v2

# Paramagnon-Interference Mechanism for Three-Dimensional Bond Order in Kagome Metals AV3Sb5 (A=Cs, Rb, K): Analysis by the Density-Wave Equation

Abstract
The mechanism of CDW and its 3D structure are important fundamental issues in kagome metals. We have previously shown that, based on a 2D model, $2\times 2$ bond order (BO) emerges due to the paramagnon-interference (PMI) mechanism and that its fluctuations lead to $s$-wave superconductivity. This paper studies these issues based on realistic 3D models of kagome metals AV3Sb5 (A=Cs, Rb, K). We reveal that a commensurate 3D $2\times 2\times 2$ BO is caused by the PMI mechanism, by performing the 3D density-wave (DW) equation analysis for all A=Cs, Rb, K models in detail. Our results indicate a BO transition temperature $T_{\rm BO}\sim 100$K within the regime of moderate electron correlation. The 3D structure of BO is attributed to the three-dimensionality of the Fermi surface, while the 3D structure of BO is sensitively changed, since the Fermi surface is quasi-2D. Based on the analysis of the DW equation, by taking into account a finite third-order Ginzburg-Landau (GL) term, (i) shift stacking $2\times 2\times 2$ BO can be realized via a first-order transition below $T_{\mathrm{BO}}$. Here, the in-plane BO pattern (tri-hexagonal or star-of-David) is determined by the sign of the third-order GL term, with hole doping tending to favor the tri-hexagonal state. On the other hand, if the third-order GL term is very small, (ii) alternating vertical stacking BO may instead be realized via a second-order transition. The present study enhances our understanding of the rich variety of BOs observed experimentally. It is confirmed that the PMI mechanism is the essential origin of the 3D CDW of kagome metals.
I Introduction
The multistage rich quantum phase transitions in the kagome-lattice superconductors AV3Sb5 (A=Cs, Rb, K) attract significant attention in the field of condensed matter physics [2, 3, 4]. In each V-Sb plane, the $\times2$ charge density wave (CDW) has been observed below $T_{\rm BO}\approx 90$ K at ambient pressure [5, 6]. The structure of CDW is tri-hexagonal (TrH) or star-of-David (SoD) caused by the triple-$\bm{q}$}} ($3Q$) bond order (BO), which is the even-parity modulation in the hopping integral $\delta t_{i,j}^{\rm b}$ (=real) [7, 8, 9, 10, 11, 12, 13, 14, 15]. Below $T_{\rm BO}$, superconductivity (SC) appears at $T_{\rm SC}\approx 1\sim 3$ K, and it increases to 8 K under $P\approx 3{\rm GPa}$ [4]. At ambient pressure for A=Cs, the highly anisotropic $s$-wave SC state is realized [16, 17], which is naturally understood based on the BO fluctuation mechanism [14].
In a single V-Sb plane, the order parameters of the $3Q$ BO are expressed as ${\bm{\phi}}=(\phi_{1},\phi_{2},\phi_{3})$, where $\phi_{m}$ is the strength of the $m$th BO at wavevector ${{\bm{q}}}_{m}$ ($m=1,2,3$) as shown in Fig. 1 (a). The TrH and SoD BOs are shown in Figs. 1 (b) and (c), respectively. Each $3Q$ order is composed of 4 unit cells of the kagome lattice, that is, 12 V atoms.
It is important to understand the BO state because it is an essential electronic property in kagome metals. Many theoretical researches have been performed. The BO is formed in the presence of strong off-site interaction $I$. Large $I$ is mediated by the effective off-site electron correlations [7, 8, 11, 12, 14]. The bond-stretching phonon also mediates $I$ [9]. The absence of acoustic phonon anomaly at $T_{\rm BO}$ [18] indicates the important contribution from the electron correlations. In Ref. [14], the authors found that the BO is naturally induced by the “paramagnon-interference (PMI) mechanism”, which is dropped in the mean-field-type approximations. This mechanism naturally explains the orbital and bond orders observed in Fe-based and cuprate superconductors [19].
Recently, the three-dimensional (3D) nature of the BO in kagome metals has attracted increasing attention because it drastically alters their electronic properties. Depending on the carrier doping and the pressure, the BO states with period-1, period-2, and period-4 along the $c$-axis, which are called the $2\times 2\times C$ BO ($C=1,2,4$), have been reported. As shown in Figs. 2 (a) and (b), in the case of $C=2$, TrH-SoD alternating vertical stacking BO (v-BO) is realized for $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,\pi\}$, while the nematic TrH or SoD shift stacking BO (s-BO) is realized for $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,0\}$, where $q_{m}^{z}$ is the $z$ component of wavevector in the $m$th BO. In the nematic s-BO, $C_{6}$ symmetry is broken to $C_{2}$. The nematic director is represented by the arrow connecting the $C_{6}$ center of the upper layer to that of the lower layer, as shown in Fig. 2 (b).
In addition, the time-reversal-symmetry-breaking TRSB state has been observed by various experimental methods [5, 20, 21, 22, 23, 24, 25, 26, 27]. As for the origin of the TRSB state, the chiral charge current order, which is the pure imaginary BO $\delta t_{i,j}^{\rm c}$ (=imaginary), has been proposed [11, 12, 13, 15, 28, 29]. This topic attracts significant attention in kagome metals. The recent theoretical study [15] has revealed that the current order structure is strongly coupled to the background BO. Thus, the 3D nature of the BO is very important for understanding nematicity and the chiral current order.
In this paper, we study the 3D nature of the density wave (DW) in kagome metals based on the realistic 3D multiorbital models for AV3Sb5 (A=Cs, Rb, K), by using the 3D DW equation method [14, 30, 31, 32]. We reveal that commensurate $2\times 2\times 2$ BO is realized by the PMI mechanism described by the Aslamazov-Larkin vertex corrections (AL-VCs) in all A=Cs, Rb, K models. The 3D structure of BO originates from that of the Fermi surface. Based on the analysis of the DW equation, by taking into account a finite third-order Ginzburg-Landau (GL) term, (i) $2\times 2\times 2$ s-BO can be realized via a first-order transition below $T_{\mathrm{BO}}$. Here, the in-plane BO pattern (TrH or SoD) is determined by the sign of the third-order GL term, with hole doping tending to favor the TrH state. On the other hand, if the third-order GL term is very small, (ii) alternating v-BO may instead be realized via a second-order transition. These results are robust for all A=Cs, Rb, and K models. The present study enhances our understanding of the rich variety of BOs observed experimentally. It is confirmed that the PMI mechanism is the essential origin of the 3D CDW of kagome metals.
Figure 1: (a) 2D wavevector $\bm{q}_{m}$ $(m=1,2,3)$ in the original 2D Brillouin zone. (b) 2D TrH BO that corresponds to ${\bm{\phi}}\propto(1,1,1)$. In the BO state, the thick purple bond $(i,j)$ corresponds to a decrease in V-V bond length. (c) 2D SoD BO that corresponds to ${\bm{\phi}}\propto(-1,1,1)$.
Figure 2: Possible $2\times 2\times 2$ 3D BO states: (a) “alternating v-BO” with $C_{6}$ symmetry. (b) “TrH s-BO” with nematic $C_{2}$ symmetry. The red arrow represents the nematic director.
Here, we present a brief summary of the CDW structure observed in AV3Sb5. In each layer, the $2\times 2$ superlattice structure that indicates the formation of the $3Q$ CDW has been observed for A=Cs, Rb, and K [5, 33, 34, 35, 36, 37, 38, 39, 40]. Recently, the $2\times 2\times 1$ superlattice [41], the $2\times 2\times 2$ superlattice [18, 41, 47, 43, 44, 45, 42, 46, 48, 49, 50, 51, 52, 53], and the $2\times 2\times 4$ superlattices [41, 51, 52, 53, 54, 55] have been observed in the CDW phase. In this study, we take into account the $2\times 2\times 1$ and $2\times 2\times 2$ CDW states for simplicity.
In the $2\times 2\times 2$ CDW state, (i) the nematic TrH or SoD s-BO has been supported by the STM [42], the NMR [43, 44, 45, 46, 37, 56], the ARPES [47, 48], the Raman [49], and the X-ray [18, 50, 51] measurements for A=Cs, Rb, and K. In general, it is difficult to distinguish between the TrH and SoD BOs in experiments such as STM and X-ray measurements, since the observed results can be interpreted by both the TrH and SoD BOs. On the other hand, (ii) the alternating v-BO has been proposed by combining the ARPES measurement and the DFT first-principles calculation in A=Cs [48]. Consensus on the BO states in kagome metals has yet to be reached, since the observed BO depends on doping, pressure, and experimental method.
Theoretically, the TrH stacking CDW due to the electron-phonon interaction mechanism is obtained by the DFT first-principles studies [9, 57, 58, 59]. However, the change in bond length obtained by the first-principles studies is rather large compared to the experimental results [41, 51, 54]. Thus, it is reasonable to assume that the origin of the CDW is the electron correlation. As for the electron correlation mechanism, the 3D CDW formation due to the on-site and off-site Coulomb interaction has been studied based on the mean-field theory [60, 61] and the variational method [62], by introducing sizable off-site interactions.
In contrast, the present authors studied the realistic 2D kagome metal model based on the DW equation method by considering only the on-site Coulomb interaction $U$ [14]. In the PMI mechanism, the effective off-site interaction is induced by the AL terms, which are the higher-order electron correlation. We revealed that the commensurate $2\times 2$ CDW is naturally and robustly given by the PMI mechanism. In addition, $s$-wave superconductivity is explained by the CDW fluctuations. In this paper, we show that the same mechanism can explain the 3D CDW based on realistic 3D kagome metal models.
GL theories in 2D kagome metals have revealed that the $3Q$ $2\times 2$ TrH or SoD CDW state is stabilized by the 3rd-order GL terms [12, 15, 61]. The GL theories also expect the coexistence of the 3Q CDW and chiral-charge-current orders. In the coexisting state, $C_{2}$ nematicity appears [15, 61].
The 3D structure of CDW has been discussed in Refs. [12, 63, 64] by using the 3D GL theory. The $2\times 2\times 2$ s-BO has been proposed by the 3D GL theory. In CsV3Sb5, the TrH s-BO was proposed [63, 64], where the coefficients of the GL terms are estimated by the first-principles calculation.
II 3D Multiorbital Model for Kagome Metal
Here, we derive a first-principles realistic tight-binding model for CsV3Sb5, by using the WIEN2k [65] and Wannier90 [66] software. Hereafter, the unit of energy is eV unless otherwise noted. The derived kinetic term is
$H_{0}=\sum_{{{\bm{k}}},lm{\sigma}}h_{lm}^{0}({{\bm{k}}})c_{{{\bm{k}}}l{\sigma}}^{\dagger}c_{{{\bm{k}}}m{\sigma}},$ (1)
where $\bm{k}$}} is the 3D wavevector, $l,m$ represents the V $d$-orbitals and Sb $p$-orbitals, and $\sigma$} is the spin index. This is a 30 orbital tight-binding model. To improve the tight-binding model, we introduce the energy shift of all Sb $p$-orbitals $\Delta E_{p}=-0.2$. The bandstructure compared to that for $\Delta E_{p}=0$ is shown in Fig. 9 (a) in Appendix A. Energy at the $\Gamma$ point $E_{\Gamma}\sim-0.4$ given for original $\Delta E_{p}=0$ is reduced to $E_{\Gamma}\sim-0.55$ by introducing $\Delta E_{p}=-0.2$. Since $E_{\Gamma}\sim-0.55$ well reproduces the ARPES measurements [48, 67], $\Delta E_{p}=-0.2$ is reasonable for the model of CsV3Sb5. In this paper, based on the first-principles calculation, the thick purple bond in Figs 1 and 2, where the V-V bond length decreases, corresponds to a modulation of the hopping integral $\delta t>0$.
The 3D FS structure is shown in Fig. 9 (b) in Appendix A. The FSs at $n=31$ (=original electron number) are shown in Figs. 3 (a)-(c). Also, the FSs at $n=30.8$ (hole doping) are shown in Figs. 3 (d)-(f). These FSs are quasi-2D. The FSs in $k_{z}=0$ are different between $n=31$ and $n=30.8$ due to the Lifshitz transition around the M point, while the FSs in $k_{z}=\pi$ at $n=31$ are almost the same as those at $n=30.8$. The FSs of A=Rb, K for $E_{p}=-0.2$ are shown in Fig. 4. The FSs of A=Rb, K are similar to those of the hole-doped A=Cs in Fig. 3. The bandstructures of A=Rb, K for $E_{p}=-0.2$ are shown in Fig. 10 in Appendix A. They are similar to that of A=Cs. $E_{\Gamma}\sim-0.6$ for A=Rb, K is also consistent with the ARPES measurements [48].
Next, we calculate the effects of the strong Coulomb interaction of $3d$-orbital electrons on V sites. In kagome metals, the FSs are mainly composed of two of the five $d$-orbitals of V atom: The “pure-type” FS is composed of three $d_{xz}$ (or $b_{3g}$ rep.) orbitals on sublattices A, B, C, and the “mix-type” FS is composed of three $d_{yz}$ orbitals (or $b_{2g}$ rep.). In the pure-type FS, the bandstructure around each van Hove singularity vHS-$S$ (SS=A, B, C) is composed of the single sublattice $S$, which is known as the “sublattice interference”. In contrast, in the mix-type FS, each vHS-$S$ is composed of two of three sublattices. Both the pure-type and mix-type bands are nearly independent, although they weakly hybridize each other.
[…]
Here, we study the spin-channel and charge-channel susceptibilities, ${\hat{\chi}}^{s}(q)$ and ${\hat{\chi}}^{c}(q)$, based on the random-phase-approximation (RPA), which is equivalent to the mean-field approximation. To perform the diagrammatic calculation, we introduce the $30\times 30$ matrix expression of the Green function as
${\hat{G}}(k)=[(i{\epsilon}_{n}+\mu){\hat{1}}-{\hat{h}}^{0}({{\bm{k}}})]^{-1},$ (2)
where $\mu$ is the chemical potential, $k\equiv({{\bm{k}}},{\epsilon}_{n})$, and ${\epsilon}_{n}=(2n+1)\pi T$ is the fermion Matsubara frequency. In the RPA, ${\hat{\chi}}^{s}(q)$ and ${\hat{\chi}}^{c}(q)$ are given as
${\hat{\chi}}^{s(c)}(Q)$ $=$ ${\hat{\chi}}^{0}(q)\left[{\hat{1}}-{\hat{\Gamma}}^{s(c)}{\hat{\chi}}^{0}(Q)\right]^{-1},$ (3)
$\chi^{0}_{ll^{\prime},mm^{\prime}}(Q)$ $=$ $-\frac{T}{N}\sum_{k}G_{lm}(k+Q)G_{m^{\prime}l^{\prime}}(k),$ (4)
where $Q\equiv({{\bm{Q}}},{\omega}_{l})$ with the bosonic Matsubara frequency ${\omega}_{l}=2\pi Tl$, and $\chi^{0}_{ll^{\prime},mm^{\prime}}(Q)$ is the irreducible susceptibility. Here, $\Gamma_{ll^{\prime}mm^{\prime}}^{s(c)}$ is the 5-orbital Coulomb interaction for the same V sublattice in the spin (charge) channel, composed of the intra-orbital $U$, inter-orbital $U^{\prime}$, and the exchange term $J=(U-U^{\prime})/2$. The expression of $\Gamma_{ll^{\prime}mm^{\prime}}^{s(c)}$ is presented in Ref. [68, 69]. We use $\bm{k}$}} meshes $N=N_{x}\times N_{y}\times N_{z}=60\times 60\times 32$, and Matsubara frequency $N_{\omega}=512$. Here, we introduce the spin (charge) Stoner factor ${\alpha}_{S(C)}$ given as the largest eigenvalue of ${\hat{\Gamma}}^{s(c)}{\hat{\chi}}^{0}(q)$. $\chi^{s(c)}$ diverges when ${\alpha}_{S(C)}$ reaches unity. In the RPA, ${\alpha}_{S}$ is always larger than ${\alpha}_{C}$ [68, 69] for $U>U^{\prime}$, so the nonmagnetic BO in kagome metals cannot be explained unless large nearest-neighbor Coulomb interaction $V$ ($V>0.5U$) exists. However, the BO and the current order are naturally explained by considering the beyond-mean-field electron correlations described by the VCs.
In previous theoretical studies [14], both the spin and BO fluctuations develop in pure-type and mix-type FSs. It was shown that the electron correlations are stronger on pure-type FS composed of $d_{xz}$ (or $b_{3g}$) orbitals. Therefore, in the present study, we consider the on-site Coulomb interactions only on the $b_{3g}$ orbitals. Here, we assign $l=1,2,3$ to the $b_{3g}$ orbital on sites A, B, C, respectively. In this case, the Coulomb interaction is simply given as ${\hat{\Gamma}}^{s(c)}_{mm^{\prime},ll^{\prime}}=U^{s(c)}\delta_{m,m^{\prime}}\delta_{l,l^{\prime}}\delta_{l,m}$, where $U^{s}=U$ and $U^{c}=-U$. Then, the RPA susceptibility is simplified as
${\hat{\tilde{\chi}}}^{s(c)}(Q)$ $=$ ${\hat{\tilde{\chi}}}^{0}(Q)\left[{\hat{1}}-(+)U{\hat{\tilde{\chi}}}^{0}(Q)\right]^{-1},$ (5)
where ${\hat{\tilde{\chi}}}^{s(c)}(q)$ is $3\times 3$ matrix: ${\tilde{\chi}}_{l,m}^{0}(Q)=\chi^{0}_{ll,mm}(Q)$ with $l,m=1,2,3$. In the next section, we introduce the DW equation to study the BO in kagome metals. The kernel function of the DW equation is composed of the RPA susceptibility.
Hereafter, we denote a 3D variable wavevector ${{\bm{Q}}}=({{\bm{q}}},q^{z})$, where $\bm{q}$}} ($q^{z}$) is the 2D (1D) variable wavevector. We also introduce the notation ${{\bm{Q}}}_{m}=({{\bm{q}}}_{m},q_{m}^{z})$ ($m=1,2,3$), where ${{\bm{q}}}_{m}$ is the fixed 2D wavevector introduced in Fig. 1 (a), while $q_{m}^{z}$ is treated as a variable.
We introduce the “form factor” of the BO at wavevector $\bm{Q}$}}, ${\hat{g}}_{{{\bm{Q}}}}(k)$, which is given by the Fourier transformation of the 3D BO, such as the 3D stacking of Figs. 1 (b) and (c). The form factor is normalized by setting ${\rm max}|{\hat{g}}_{{{\bm{Q}}}}(k)|=1$. Here, ${\hat{g}}_{{{\bm{Q}}}}(k)$ is a $3\times 3$ matrix with respect to the three sublattices A, B, and C. In this study, we derive the BO form factor and its eigenvalue $\lambda_{{\bm{Q}}}$ in 3D momentum space microscopically, by analyzing the DW equation for the 3D kagome metal models. It is found that both ${\hat{g}}_{{{\bm{Q}}}_{m}}(k)$ and $\lambda_{{{\bm{Q}}}_{m}}$ ($m=1,2,3$) have very small $q_{m}^{z}$-dependences.
III DW Equation Analysis for 3D Kagome Metal Model
Here, we study the BO in the 3D kagome metal models based on the DW equation theory. In this theory, various nonmagnetic DW orders are caused by the beyond-RPA non-local correlations. This theory was first developed to explain the orbital nematic state in Fe-based superconductors, and it has been successfully applied to the cuprates [70] and the twisted-bilayer graphene [71]. Recently, the present authors studied 2D kagome metals in Ref. [14]. By solving the DW equation including the Hartree, Maki-Thompson (MT), and AL terms shown in Fig. 5 (a), we can calculate the infinite series of AL and MT terms. The AL terms represent the PMI mechanism, which gives rise to the BO in kagome metals. The analytic expression of the charge-channel DW equation is
$\lambda_{{{\bm{Q}}}}g_{{\bm{Q}}}^{L}(k)$ $=$ $-\frac{T}{N}\sum_{p,M_{1},M_{2}}I_{{\bm{Q}}}^{L,M_{1}}(k,p)$ (6)
$\times\{G(p)G(p+{{\bm{Q}}})\}^{M_{1},M_{2}}g_{{\bm{Q}}}^{M_{2}}(p),$
where $k\equiv({{\bm{k}}},{\epsilon}_{n})$ and $p\equiv({{\bm{p}}},{\epsilon}_{m})$. The largest eigenvalue $\lambda_{{\bm{Q}}}$ represents the DW instability at wavevector $\bm{Q}$}}, and the corresponding form factor ${\hat{g}}_{{\bm{Q}}}$ gives the DW order function. $L\equiv(l,l^{\prime})$ and $M_{i}$ represent the pair of sublattice indices $A,B,C$. The phase factor of the eigenvector is fixed to satisfy the Hermitian condition $g_{{\bm{Q}}}^{ll^{\prime}}({{\bm{k}}},{\epsilon}_{n})=[g_{-{{\bm{Q}}}}^{l^{\prime}l}({{\bm{k}}}+{{\bm{Q}}},-{\epsilon}_{n})]^{*}$. In addition, we fix the sign of ${\hat{g}}_{{{\bm{Q}}}_{m}}$ so that $\phi_{m}{\hat{g}}_{{{\bm{Q}}}_{m}}$ reproduces the BO shown in Fig. 1.
The effective interaction ${\hat{I}}_{{\bm{Q}}}(k,p)$ in the DW equation is given as
$I^{ll^{\prime},mm^{\prime}}_{{{\bm{Q}}}}(k,k^{\prime})=\sum_{b=s,c}\frac{a^{b}}{2}\Bigl[-V^{b}_{lm,l^{\prime}m^{\prime}}(k-k^{\prime})$
$+\frac{T}{N}\sum_{p}\sum_{l_{1}l_{2},m_{1}m_{2}}V^{b}_{ll_{1},mm_{1}}\left(p+{{\bm{Q}}}\right)V^{b}_{m^{\prime}m_{2},l^{\prime}l_{2}}\left(p\right)$
$\qquad\qquad\qquad\quad\times G_{l_{1}l_{2}}(k-p)G_{m_{2}m_{1}}(k^{\prime}-p)$
$+\frac{T}{N}\sum_{p}\sum_{l_{1}l_{2},m_{1}m_{2}}V^{b}_{ll_{1},m_{2}m^{\prime}}\left(p+{{\bm{Q}}}\right)V^{b}_{m_{1}m,l^{\prime}l_{2}}\left(p\right)$
$\qquad\qquad\qquad\times G_{l_{1}l_{2}}(k-p)G_{m_{2}m_{1}}(k^{\prime}+p+{{\bm{Q}}})\Bigr],$ (7)
where $a^{s(c)}=3$(11) and $p=({{\bm{p}}},{\omega}_{l})$. $\hat{V}^{b}$ is the $b$-channel interaction given by $\hat{V}^{b}=\hat{\Gamma}^{b}+\hat{\Gamma}^{b}\hat{\chi}^{b}\hat{\Gamma}^{b}$.
Figure 5: (a) Charge-channel density-wave (DW) eigenvalue equation. The largest eigenvalue $\lambda_{{\bm{Q}}}$ represents the instability of the charge-channel DW at wavevector $\bm{Q}$}}, and the form factor $f$ gives the order parameter. The wavy lines represent the spin susceptibilities. Two AL terms describe the paramagnon interference mechanism that gives various unconventional DW states. (b-c) Obtained $\lambda_{{\bm{Q}}}$ for A=Cs at $n=30.8,31$: (b) $\lambda_{{\bm{Q}}}$ in the $q^{z}=0$ plane, where ${{\bm{q}}}_{K}=\left(\frac{2\pi}{\sqrt{3}},\frac{2\pi}{3}\right)$. (c) Small $q_{1}^{z}$-dependence of $\lambda_{{{\bm{Q}}}_{1}}$ for ${{\bm{Q}}}_{1}=({{\bm{q}}}_{1},q_{1}^{z})$. The BO solution is obtained for any $q_{1}^{z}$. (d) $\lambda_{{{\bm{Q}}}}$ for $\alpha_{S}=0.95,0.98$ in $q^{z}=0$ at $n=31$ of A=Cs. (e) $\alpha_{S}$ dependence of $\lambda_{{{\bm{Q}}}_{1}}$ for $q_{1}^{z}=0$ at $n=31$ of A=Cs.
Note that the form factor is proportional to the particle-hole condensation $\sum_{\sigma}\langle c_{{{\bm{k}}}+{{\bm{q}}},l,{\sigma}}^{\dagger}c_{{{\bm{k}}},l^{\prime},{\sigma}}\rangle$, or equivalently, the symmetry-breaking component in the self-energy.
Here, we explain the obtained numerical results for A=Cs at $T=0.04$ and $U=2.6$, where ${\alpha}_{S}=0.90$. Figure 5 (b) shows the obtained $\lambda_{{\bm{Q}}}$ on the $q^{z}=0$ plane at $n=30.8,31$: $\lambda_{{\bm{Q}}}$ shows the almost commensurate peak at ${{\bm{q}}}={{\bm{q}}}_{1}$, which corresponds to the 2D BO wavevector in kagome metals. Interestingly, the $q_{1}^{z}$-dependence of $\lambda_{{{\bm{Q}}}_{1}}$ is quite small ($\lesssim 1$%), as shown in Fig. 5 (c), reflecting the quasi-2D character of the FSs in kagome metals. We note that the value of $\lambda_{{{\bm{Q}}}_{1}}$ for $q_{1}^{z}=\pi$ is slightly larger than that for $q_{1}^{z}=0$ since the spin susceptibility peaks in $q^{z}\sim\pi$ due to the 3D structure of FSs, which derives the wavevector with $q^{z}\sim\pi$ as discussed in Ref. [31]. In this case, the alternating v-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,\pi\}$ is the most favorable, while the TrH (SoD) s-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,0\}$ is also favored. Within the DW equation, we cannot distinguish between the TrH s-BO and the SoD s-BO. Figures 5 (d) and (e) show $\alpha_{S}$ dependence of $\lambda_{{{\bm{Q}}}}$ for $q^{z}=0$ at $n=31$, $T=0.04$. $\lambda_{{{\bm{Q}}}_{1}}$ reaches $1$ for $\alpha_{S}\gtrsim 0.97$. Thus, the BO transition temperature $T_{\rm BO}\sim 100$K is obtained in the moderate electron correlation.
Figure 6: (a-b) Obtained $\lambda_{{\bm{Q}}}$ for A=Rb, K at $n=31$: (a) $\lambda_{{\bm{Q}}}$ in the $q^{z}=0$ plane. (b) Small $q_{1}^{z}$-dependence of $\lambda_{{{\bm{Q}}}_{1}}$ for ${{\bm{Q}}}_{1}=({{\bm{q}}}_{1},q_{1}^{z})$.
In deriving Figs. 5 (b)-(e), we take only the Coulomb interaction on the $b_{3g}$ orbital into account. This simplification does not alter the qualitative results, while the quantitative results are improved by considering the multiorbital Coulomb interaction.
Figures 6 (a) and (b) show $\bm{Q}$}} dependence of $\lambda_{{\bm{Q}}}$ for A=Rb, K at $T=0.04$ with ${\alpha}_{S}=0.90$ for $U=2.5$. The obtained results are similar to those for A=Cs shown in Fig. 5. Thus, it is robust that the alternating v-BO and the TrH (SoD) s-BO are favored in kagome metals.
Figure 7: (a) Obtained form factor $g_{{{\bm{Q}}}_{1}}^{\rm AB}({{\bm{k}}})$ in the $k_{x}$-$k_{y}$ plane at ${{\bm{q}}}_{1}^{z}=0$ at $n=31$. (b) $\bm{k}$}} dependence of $g_{{{\bm{Q}}}_{1}}^{\rm AB}({{\bm{k}}})$ for $q_{1}^{z}=0,\pi$ along the blue arrow in (a).
Figure 7 (a) shows the obtained form factor $g_{{{\bm{Q}}}_{1}}^{\rm AB}({{\bm{k}}})$ in the $k_{x}$-$k_{y}$ plane at ${{\bm{q}}}_{1}^{z}=0$ in the case of $n=31$. Here, $\hat{g}_{{{\bm{Q}}}}({{\bm{k}}})\equiv\hat{g}_{{{\bm{Q}}}}({{\bm{k}}},0)$ is given by the analytic continuation (linear extrapolation) of $\hat{g}_{{{\bm{Q}}}}(k)$. We also show the $\bm{k}$}} dependence of $g_{{{\bm{Q}}}_{1}}^{\rm AB}({{\bm{k}}})$ for $q_{1}^{z}=0,\pi$ at $k_{z}=0$ in Fig. 7 (b). $q_{1}^{z}$ dependences of $g_{{{\bm{Q}}}_{1}}^{\rm AB}({{\bm{k}}})$ are quite small. We confirm that $k_{z}$ dependences of $g_{{{\bm{Q}}}_{1}}^{\rm AB}({{\bm{k}}})$ is also negligible. Thus, we ignore the $q^{z}$ and $k_{z}$ dependences of the form factors.
To summarize, the eigenvalue for the BO $\lambda_{{\bm{Q}}}$ derived from the DW equation analysis for the 3D multiorbital kagome metal models for A=Cs, Rb, K shows the maximum value at commensurate ${{\bm{Q}}}={{\bm{Q}}}_{m}$ ($m=1,2,3$), while its $q_{m}^{z}$ dependence is small. This result would be consistent with the rich variety of the 3D BO states observed in kagome metals. In the present study, $\lambda_{({{\bm{q}}}_{m},\pi)}$ is slightly larger than $\lambda_{({{\bm{q}}}_{m},0)}$. In this case, the alternating v-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,\pi\}$ in Fig. 2 (a) is the most favorable, while the TrH (SoD) s-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,0\}$ in Fig. 2 (b) is also favored. These results are robust for all A=Cs, Rb, and K models. The rich variety of BOs observed in experiments would be understood by the competition between the alternating v-BO and TrH (SoD) s-BO. We verify that the PMI mechanism in the DW equation is the essential origin of the 3D CDW in kagome metals.
IV Discussion
Here, we discuss the 3D structure of CDW in kagome metal based on the obtained results by the 3D DW equation and the 3rd-order GL free energy. For simplicity, we take account of commensurate $q_{m}^{z}=-q_{m}^{z}(=0,\pi)$ BO wavevectors, which correspond to $2\times 2\times 1$ and $2\times 2\times 2$ BO. From the $\bm{Q}$}} dependence of $\lambda_{{\bm{Q}}}$ shown in Figs. 5 and 6, we find that $\lambda_{{{\bm{q}}}_{1},\pi}$ is slightly larger than $\lambda_{{{\bm{q}}}_{1},0}$. For this reason, the commensurate $2\times 2\times 2$ alternating v-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,\pi\}$ and nematic TrH (SoD) BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,0\}$ are favored based on the 3D DW equation due to the PMI mechanism. On the other hand, the coefficient of the 3rd-order GL term $b_{1}(q_{1}^{z},q_{2}^{z},q_{3}^{z})$ is shown in Fig. 11 (b) in Appendix B. The obtained negative (positive) $b_{1}$ favors the TrH (SoD) BO. We confirmed that $b_{1}(\pi,\pi,0)$ is almost the same as $b_{1}(0,0,0)$ since $q^{z}$ dependence of $b_{1}$ is very small. Thus, $2\times 2\times 2$ TrH (SoD) s-BO competes with $2\times 2\times 1$ TrH (SoD) BO. Combined with the results of the 3D DW equation, the nematic $2\times 2\times 2$ TrH (SoD) s-BO is realized via a first-order transition below $T_{\mathrm{BO}}$ by the 3rd-order GL free energy. In contrast, the contribution from the 3rd-order GL free energy is absent to the alternating v-BO with $\{q_{1}^{z},q_{2}^{z},q_{3}^{z}\}=\{\pi,\pi,\pi\}$, which violates the constraint $q_{1}^{z}+q_{2}^{z}+q_{3}^{z}=0$ (modulo $2\pi$). The alternating v-BO will be realized via a second-order transition if the 3rd-order GL free energy is very small. Based on the $n$ dependence of $b_{1}$, the TrH BO tends to be stabilized in the hole-doped regime, whereas the SoD BO tends to be favored in the electron-doped regime. However, the obtained value of $b_{1}$ is sensitive to $T$. Quantitative evaluation of $b_{1}$ is an important future issue.
In the NMR study for A=Cs [37, 56], the SoD BO has been proposed at low temperatures, while other NMR studies have proposed the TrH s-BO near the CDW transition temperature. Since the value of $b_{1}$ at $n=31$ is not robust, as shown in Figs. 11 (b), the TrH and SoD BOs can compete. These BOs are favored by the DW equation due to the PMI mechanism, where the alternating v-BO is also favored. These competitions would explain the rich variety of the 3D DW states in kagome metals.
Finally, we clarify the relation between a modulation of the V-V length $\delta a$ and a BO parameter $\delta t$. We compare the band structure obtained from a first-principles calculation using $\delta a$ with that obtained from a model calculation using $\delta t$. In the following, we set $\Delta E_{p}=0$. The relation between $\delta a$ and $\delta t$ is essentially unaffected by the value of $\Delta E_{p}$.
Figures 8 (a) and (b) present the 2D folded $2\times 2$ band structures for A=Cs obtained from the first-principles calculations using $\delta a=0$ and the TrH-type modulation $\delta a/a_{0}=0.01$, respectively. $a_{0}$ denotes the original V-V length without modulation. Hereafter, we focus on the $d_{xz}$ orbital, the weight of which is represented by color intensity in these figures. The $2\times 2$ TrH-type modulation shown in Fig. 1 (b) is characterized by a shortened bond length $a_{0}-\delta a$ on the thick purple bond and an elongated bond length $a_{0}+\delta a$ on other bonds. Due to the $2\times 2$ band folding, three vHSs at M points in the original Brillouin zone are folded to $\Gamma$ point. In the $2\times 2$ TrH-type modulation, the triply degenerate bands at $\Gamma$ point split into one upper and two lower bands. This band splitting is also observed for A = K, as shown in Figs. 8 (c) and (d). The main difference between the results for A=Cs and A=K is the energy of vHS $E_{\rm vHS}$; $E_{\rm vHS}\sim-0.05$eV for A=K is higher than $E_{\rm vHS}\sim-0.1$eV for A=Cs.
Here, we compare the band splitting obtained by the first-principles calculation using the TrH-type modulation with that obtained by a model calculation using the BO parameter $\delta t$. The band splittings by the TrH-type modulation in Figs. 8 (b) and (d) are effectively reproduced using a simple model with a positive $\delta t>0$, as shown in Figs. 8 (e) and (f). In the simple model introduced in Ref. [14], only the three $d_{xz}$ orbitals are taken into account. In this model, the inter-V hopping is given as $t_{0}+\delta t$ on the thick purple bond, while that on other bonds is $t_{0}-\delta t$, where $t_{0}<0$ denotes the original hopping, and the BO parameter is $\delta t>0$. Interestingly, the magnitude of the inter-V hopping $|t_{0}+\delta t|$ decreases as the V-V bond length shortens. This counterintuitive behavior arises from the competition between direct V-V hopping $t_{\rm direct}$ and indirect V-Sb-V hopping $t_{\rm indirect}$. The inter-V hopping is divided into $t_{\rm direct}$ and $t_{\rm indirect}$. In the original hopping, we have $t_{0}=t_{\rm direct}+t_{\rm indirect}<0$. In this case, we confirm the relations $t_{\rm direct}>0$, $t_{\rm indirect}<0$, and $|t_{\rm direct}|<|t_{\rm indirect}|$. In general, $|t_{\rm direct}|$ and $|t_{\rm indirect}|$ exhibit different dependences on the V–V bond length. In the kagome metals, $|t_{0}+\delta t|$ decreases $(\delta t>0)$ when the V-V bond length shortens, due to the competition between $t_{\rm direct}$ and $t_{\rm indirect}$. Similar results are known for Fe-based superconductors, where the direct Fe-Fe hopping $t_{\rm direct}>0$ competes with the indirect Fe-As-Fe hopping $t_{\rm indirect}<0$ [72]. The competition between $t_{\rm direct}$ and $t_{\rm indirect}$ depends on the Fe-As-Fe angle.
Figure 8: 2D folded $2\times 2$ bandstructures of CsV3Sb5 obtained by the first-principles calculation: (a) without the V-V length modulation $\delta a=0$, and (b) with TrH-type modulation $\delta a/a_{0}=1$%. The weight of the $d_{xz}$ orbital is represented by the color intensity. Panels (c) and (d) present analogous results for KV3Sb5, corresponding to those shown for CsV3Sb5 in panels (a) and (b). (e) 2D folded $2\times 2$ bandstructure given by the simple model without BO, and (f) that with TrH BO. $\delta t>0$ denotes the BO parameter on the thick purple bond for the TrH BO.
V Summary
In this paper, we studied the 3D nature of the DW in kagome metals based on the 3D multiorbital models for AV3Sb5 (A=Cs, Rb, K), by using the DW equation method. The TrH or SoD shape commensurate BO is robustly induced in each V-Sb plane by the PMI mechanism that is described by the AL-VCs. $T_{\rm BO}\sim 100$K is obtained in the moderate electron correlation. Based on the analysis of the DW equation, by taking into account a finite third-order GL term, (i) $2\times 2\times 2$ s-BO can be realized via a first-order transition below $T_{\mathrm{BO}}$. Here, the in-plane BO pattern (TrH or SoD) is determined by the sign of the third-order GL term, with hole doping tending to favor the TrH state. On the other hand, if the third-order GL term is very small, (ii) alternating v-BO may instead be realized via a second-order transition. The rich variety of BOs observed in experiments can be understood by the combination of the DW equation analysis and the 3rd-order GL free energy. It was confirmed that the PMI mechanism is the essential origin of the 3D CDW of kagome metals.
Acknowledgements.
This work was supported by JSPS KAKENHI Grant Numbers JP25H01246, JP25H01248, JP24K00568, JP24K06938, JP23K03299, JP22K14003.
Appendix A 3D Fermi Surface and Bandstructure
We derive the 30 orbital tight-binding model with 15 V $d$-orbitals and 15 Sb $p$-orbitals by using the WIEN2k [65] and Wannier90 [66] software. Figure 9 (a) shows the bandstructures for $\Delta E_{p}=-0.2,0$. The bandstructure around the $\Gamma$ point is composed of Sb $p$-orbitals. By introducing $\Delta E_{p}=-0.2$, energy at the $\Gamma$ point $E_{\Gamma}$ decreases from 0.4-0.4eV to 0.55-0.55eV consistently with ARPES measurements [48, 67]. Figure 9 (b) shows the 3D FS for $\Delta E_{p}=-0.2$ at $n=31$. Figures 10 (a) and (b) show the bandstructures of A=Rb and A=K for $\Delta E_{p}=-0.2$, respectively. In these cases, $E_{\Gamma}\sim-0.6$ is also consistent with ARPES measurements [48].
Figure 9: (a) Bandstructure of A=Cs for $\Delta E_{P}=0$ (black line) and that for $\Delta E_{P}=-0.2$ (red line). (b) Three-dimensional FSs of A=Cs for $\Delta E_{P}=-0.2$.
Figure 10: (a) Bandstructure of A=Rb for $\Delta E_{P}=-0.2$ and (b) that of A=K.
Appendix B 3rd-Order GL Coefficient
To understand the 3D $3Q$ BO state, we also study the 3rd-order GL free energy $F^{(3)}=b_{1}(q_{1}^{z},q_{2}^{z},q_{3}^{z})\phi_{1}\phi_{2}\phi_{3}$ in the 3D kagome metal models. ${\bm{\phi}}=(\phi_{1},\phi_{2},\phi_{3})$ is the order parameters, and the potential of the BO state at ${{\bm{Q}}}={{\bm{Q}}}_{m}$ is given as $\phi_{m}{\hat{g}}_{{{\bm{Q}}}_{m}}(k)$. Here, we fix the 2D BO wavevectors ${{\bm{q}}}_{m}$ while their $q_{z}$ components $q_{m}^{z}$ are variables, under the constraint $q_{1}^{z}+q_{2}^{z}+q_{3}^{z}=0$ (modulo $2\pi$). Here, we take account of commensurate $q_{m}^{z}=-q_{m}^{z}(=0,\pi)$ BO wavevectors. Diagrammatic expression of $b_{1}(q_{1}^{z},q_{2}^{z},q_{3}^{z})$ is given in Fig. 11 (a), whose analytical expression is introduced in Ref. [15].
Figure 11 (b) shows the numerical results of $b_{1}(0,0,0)$ at $T=0.06$, $0.08$, and $0.1$ for A=Cs. The obtained negative (positive) $b_{1}$ favors the TrH (SoD) BO. We confirmed that $b_{1}(\pi,\pi,0)$ is almost the same as $b_{1}(0,0,0)$ since $q^{z}$ dependence of $b_{1}$ is very small. Thus, $2\times 2\times 2$ TrH (SoD) s-BO competes with $2\times 2\times 1$ TrH (SoD) BO. From the $n$ dependence of $b_{1}$, the TrH BO tends to be favored in the hole-doped regime, while the SoD BO tends to be favored in the electron-doped regime.
We also confirmed that the $q^{z}$ dependence of the 4th-order GL coefficients is smaller than that of the 3rd-order ones. Thus, the 3D structure of CDW would be determined by the $q^{z}$-dependence of $\lambda_{{\bm{Q}}}$ and $b_{1}$.
Figure 11: (a) Diagrammatic expression of the 3rd-order GL coefficient $b_{1}(q_{1}^{z},q_{2}^{z},q_{3}^{z})$. (b) $b_{1}(0,0,0)$ as functions of $n$ at $T=0.06$, $0.08$, and $0.1$ for A=Cs.